Find the structure using NMRC10H14 Proton NMR 7 1 PPM 58.0 19.6 integral values 46.6 -22.5 128.2 125.8 - 129.2 45.7 70E - - TT- TTTTTTTTTTTTTTTT 120 140
Answer #1

C10H14 DBE = C + 1 - H/2 = 10 + 1 - 14/2 = 4 This means there are 4 double bonds or a benzene ring. Presence of a multiplet a